Henry Inc's asset beta is 0.85. Rm = 8.5%, Rf = 2.5%, Tc = 30%. The company's target debt to equity ratio is 70%. The company's only debt is one outstanding bond issue that matures in 40 years. This bond issue has an annual coupon rate of 6%. Coupons are paid semi-annually. The bonds currently trade at 96% of face value. What is the company's WACC?
Answer #1

(a)-After-Tax Cost of Debt

  • The Yield to maturity (YTM) of the Bond is the discount rate at which the Bond’s price equals to the present value of the coupon payments plus the present value of the Face Value/Par Value
  • The Yield to maturity of (YTM) of the Bond is the estimated annual rate of return expected by the bondholders for the bond assuming that the they hold the Bonds until it’s maturity period/date.
  • The Yield to maturity of (YTM) of the Bond is calculated using financial calculator as follows (Normally, the YTM is calculated either using EXCEL Functions or by using Financial Calculator)

Variables

Financial Calculator Keys

Figure

Par Value/Face Value of the Bond [$1,000]

FV

1,000

Coupon Amount [$1,000 x 6.00% x ½]

PMT

30

Market Interest Rate or Yield to maturity on the Bond

1/Y

?

Maturity Period/Time to Maturity [40 Years x 2]

N

80

Bond Price/Current Market Price of the Bond [-$960]

PV

-960

We need to set the above figures into the financial calculator to find out the Yield to Maturity of the Bond. After entering the above keys in the financial calculator, we get the semi-annual yield to maturity on the bond (1/Y) = 3.135%.

The semi-annual Yield to maturity = 3.135%.

Therefore, the annual Yield to Maturity of the Bond = 6.27% [3.135% x 2]

The firm’s after-tax cost of debt on the Bond is the after-tax Yield to maturity (YTM)

The After-tax cost of debt = Annual Yield to maturity on the bond x (1 – Tax Rate)

= 6.27% x (1 – 0.30)

= 6.27% x 0.70

= 4.39%

(b)-The Cost of Equity

As per Capital Asset Pricing Model [CAPM], the cost of equity is calculated by using the following equation

Cost of equity = Risk-free Rate + Beta[Market rate of return - Risk-free Rate]

= Rf + B[Rm – Rf]

= 2.50% + 0.85[8.50% - 2.50%]

= 2.50% + [0.85 x 6.00%]

= 2.50% + 5.10%

= 7.60%

(c)-The company’s weighted average cost of capital (WACC)

Weighted Average Cost of Capital (WACC) = [After Tax Cost of Debt x Weight of Debt] + [Cost of equity x Weight of Equity]

= [(4.39% x (0.70/1.70)] + [7.60% x (1.00 / 1.70)]

= [4.39% x 0.4118] + [7.60% x 0.5882]

= 1.81% + 4.47%

= 6.28%

“Hence, the company’s weighted average cost of capital (WACC) will be 6.28%”

NOTE

Weight of Debt = Debt-to-equity ratio / (1 + Debt-to-equity ratio)

Weight of Equity = 1 / (1 + Debt-to-equity ratio)